OpenStudy (anonymous):

What is the n th term of the series: 1+1+2+3+5+8+13+21+...

OpenStudy (anonymous):

A(n) = A(n-1) + A(n-2) where a1 = 1 and a2 = 1 ?

OpenStudy (anonymous):

In terms of n

OpenStudy (anonymous):

Looks familiar....

OpenStudy (anonymous):

N th term is given by the formulae A(n)=A(n+1)+A(n+2) N=(n+1)+(n+2)

OpenStudy (anonymous):

???

OpenStudy (anonymous):

I know 

OpenStudy (anonymous):

But what in terms of n only.

OpenStudy (anonymous):

it's funny that there is a question about the golden ratio now

OpenStudy (anonymous):

it is fibonacci...

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

In terms of n ----> FIBOnnACCI

OpenStudy (anonymous):

\[F_n-F_{n-1}-F_{n-2}=0 \ \ \ n\ge2\] setting up characterestic equation gives\[\lambda^2-\lambda-1=0\]wchich gives\[\phi_1=\frac{1+\sqrt{5}}{2}\]\[\phi_2=\frac{1-\sqrt{5}}{2}\]and so\[F_n=A\phi_1^n+B\phi_2^n\]and all u need is finding A and B using the values of \(F_0\) and \(F_1\)

OpenStudy (anonymous):

finally

OpenStudy (anonymous):

@mukushla how